OpenStudy (anonymous):

Solve the proportion to determine the value of x. 5x−4/4x+7 = 13/11

OpenStudy (decentnabeel):

\[\frac{5x-4}{4x+11}=\frac{13}{11}\quad :\quad x=\frac{187}{3}\]

OpenStudy (anonymous):

How did you get that

OpenStudy (decentnabeel):

\[\frac{5x-4}{4x+11}=\frac{13}{11}\] \[\mathrm{Multiply\:by\:LCD=}11\left(4x+11\right)\] \[11\left(5x-4\right)=13\left(4x+11\right)\] \[\mathrm{Expand}\:11\left(5x-4\right):\quad 55x-44\] \[\mathrm{Expand}\:13\left(4x+11\right):\quad 52x+143\] \[55x-44=52x+143\] add 44 on both sides \[55x=52x+187\] subtract 52 on both sides \[3x=187\] divided 3 on both sides x=187/3

OpenStudy (decentnabeel):

are you understand @JayWasHere

OpenStudy (anonymous):

What is LCD

OpenStudy (decentnabeel):

Least Common Denominator fractions have "Common Denominators" when their denominators are the same. the "Least Common Denominator" is the smallest number that can be used for all denominators of 2 or more fractions. It makes it easy to add and subtract the fractions